11.3 An ostomy clinic uses barrier rings on its patients to help prevent ostomy-related leaks. Last year, the clinic used 6,200 units of barrier rings. Each unit costs $50.94 and the annual holding cost per unit is $10. It costs $3.50 to order each unit, and lead time for delivery is five days.

a. What is the economic order quantity for barrier rings?

b. What is the length of an order cycle?

c. Calculate the total weekly costs.

d. What is the investment cost for barrier rings?

e. When should the clinic's supply manager place an order for barrier rings, assuming no safety stocks are kept?

f. Determine the sensitivity of the answers to parts (a) through (e) to a +/–10 percent fluctuation in demand.
Answer #1

Annual demand=6200

Ordering cost=$3.50

Holding cost=$10 per unit per year

a.Optimal order quantity also known as Economic order quantity is given by the formula ,

EOQ=sqrt(2*Annual demand*cost per order/holding cost per unit per year)

=sqrt(2*6200*3.50/10)

=65.878=66 units per order

b.Length of an order cycle=EOQ/Annual demand=66/6200 =0.01064 =0.011

c.Number of orders =Annual demand/EOQ=6200/66=93.93 =94

Total Annual costs =Total holding costs +Total ordering costs

=Average inventory *holding cost per unit per year +Number of orders *cost per order

=66/2 *10 +94*3.50

=330+94*3.50

=659

Total weekly costs =Total annual costs/Number of weeks =659/52 =$ 12.67

d.Investment cost for barrier rings =Annual demand *Purchase cost =6200*50.94=$ 315828